# Bound of Casimir Effect by Holography

### v2

@@ Abstract
Inspired by the Kovtun-Son-Starinet bound, we propose that holography impose a lower bound on the Casimir effect. For simplicity, we focus on the Casimir effect between parallel planes for 3d conformal field theories and briefly comment on the generalizations to other boundary shapes and higher dimensions. Remarkably, the ghost-free holographic models impose a universal lower bound of the Casimir effect. We verify the holographic bound by free theories, Ising model, and $O(N)$ model with $N=2,3$ at critical points. Remarkably, the holographic bound is also obeyed by a general class of quantum field theories without conformal symmetries. It is interesting to find a field-theoretical proof or counterexample for the holographic bound of Casimir effect.
Introduction — Casimir effect Casimir:1948dh is a novel quantum effect that originates from the changes of vacuum fluctuations due to the boundary Plunien:1986ca ; Bordag:2001qi ; Milton:2004ya ; Bordag:2009zz . It is a candidate of dark energy Wang:2016och and plays a vital role in QCD Vepstas:1984sw and worm-hole physics Morris:1988tu ; Maldacena:2020sxe . It has been measured in experiments Mohideen:1998iz ; Bressi:2002fr ; Klimchitskaya:2009cw and has significant potential applications in nanotechnology. As the vacuum energy, Casimir energy is defined as the lowest energy of a quantum system with a boundary. Since energy should be bounded from the below, it is expected that there is a fundamental lower bound of the Casimir effect. This letter explores this vital problem. Holography (AdS/CFT) Maldacena:1997re provides a powerful tool to study the strongly coupled conformal field theory (CFT) and usually imposes bounds on various physical phenomena. The most famous example is the Kovtun-Son-Starinet (KSS) bound Kovtun:2004de for hydromechanics, which conjectures that the ratio of shear viscosity $\eta$ to entropy density $s$ has a lower bound
@@
where $\hbar$ is the reduced Planck constant and $k_{B}$ is the Boltzmann constant. The lower bound is obtained from the strongly-coupled CFT dual to Einstein gravity Policastro:2001yc and can be modified slightly by higher derivative gravity Brigante:2007nu ; Kats:2007mq ; Brigante:2008gz .
Now, let us explore whether AdS/CFT set a similar lower bound for the Casimir effect. For simplicity, we focus on the Casimir effect for 3d CFTs between parallel planes (strip) and comment on various generalizations at the end of the letter. We consider flat space and impose the same boundary conditions on the two planes. Then, the Casimir effect takes a universal form
@@
where $\kappa_{1}$ is the dimensionless Casimir amplitude, $c$ is the speed of light, $L$ is the distance between the two plates. For a CFT with fixed strip width $L$, $(-\kappa_{1})$ should be bounded from below to avoid negative infinity energy density $(T_{tt}=-\kappa_{1}\hbar c/L^{3})$ and pressure $(T_{nn}=-2\kappa_{1}\hbar c/L^{3})$. Note that the Casimir amplitude $\kappa_{1}$ is an extensive parameter in the sense that if one CFT has a Casimir amplitude $\kappa_{1}$, then $N$-copies of the same CFT has a Casimir amplitude $N\kappa_{1}$. It implies only $(-\kappa_{1})$ over another extensive parameter could have a lower bound. Natural candidates of these extensive parameters are the central charges of CFTs. Since the Casimir effect is a boundary effect, the most natural candidates are boundary central charges defined in Weyl anomalyJensen:2015swa $\mathcal{A}=\int_{\partial M}d^{2}x\sqrt{|\sigma|}\Big{(}aR_{\partial M}+b\text{tr}\bar{k}^{2}\Big{)},$ where $R_{\partial M}$ and $\bar{k}$ are the intrinsic Ricci scalar and traceless parts of extrinsic curvatures on the boundary $\partial M$, respectively. $a$ is the A-type boundary central charge, which decreases under the boundary RG flow Jensen:2015swa . We remark that $a$ can be positive, zero, or negative. For example, $a=\pm 1/(384\pi)$ for conformally coupled free scalar with Robin boundary condition (RBC) and Dirichlet boundary condition (DBC) Jensen:2015swa , respectively; $a\sim\sinh(\rho)$ with $-\infty<\rho<\infty$ for holographic CFT3 Fujita:2011fp . Since $a$ can be zero and negative, there is no lower bound for $(-\kappa_{1}/a)$. On the other hand, the B-type boundary central charge $b$ is always positive since it is related to the Zamolodchikov norm of displacement operator, i.e., $b=C_{D}\pi/32>0$. The displacement operator $D(y)$ describes the violation of space translation invariance normal to the boundary Billo:2016cpy , i.e., $\nabla_{i}T^{ij}=\delta(x)D(y)n^{j}$. The two-point function defines its Zamolodchikov norm, i.e., $\langle D(y)D(0)\rangle=C_{D}/|y|^{6}$. By definition, $C_{D}=32b/\pi$ is always positive. Thus, we propose the ratio of Casimir amplitude to norm of displacement operator has a lower bound set by holography
@@
Remarkably, unlike the KSS bound (1), different holographic models give the same universal lower bound (3) for Casimir effect. We verify the lower bound (3) by free CFTs, Ising model, and $O(N)$ model with $N=2,3$ at critical points. See Table. 1. Below we discuss the derivation of the holographic lower bound and its various tests and generalizations. We focus on natural units with $c=\hbar=k_{B}=1$ in the followings.
Table 1: $(-\kappa_{1}/C_{D})$ for various 3d CFTs
Fermion Scalar Ising O(2) O(3) Holography
$-1.42$ $-1.89$2.12 -1.83 -1.62 $-2.17$
Holographic bound — Let us study the holographic bound of the Casimir effect of a strip in AdS/BCFT Takayanagi:2011zk . We are interested in the general ghost-free gravity in AdS4, which turns out to be Dvali-Gabadadze-Porrati (DGP) gravity Dvali:2000hr $I=\int_{N}d^{4}x\sqrt{|g|}\Big{(}R+6\Big{)}+2\int_{Q}d^{3}y\sqrt{|h|}\Big{(}K-T+\lambda\mathcal{R}\Big{)}$, where $K$ is the extrinsic curvature, $T$ is the brane tension, $\lambda$ is the DGP parameter, $R$ and $\mathcal{R}$ are Ricci scalars in bulk $N$ and on the brane $Q$ respectively. We have set Newton’s constant $116\pi G_{N}=$ 1 and AdS radius $l=1$ for simplicity. Note that higher derivative gravity and negative DGP gravity generally suffer the ghost problems Miao:2023mui , while Gauss-Bonnet/Lovelock gravity appears only in dimensions higher than four. Thus, positive DGP gravity with $\lambda\geq 0$ is the general ghost-free gravity theory in AdS4. Following Takayanagi:2011zk , we impose Neumann boundary condition (NBC) on the brane $Q$
$K^{ij}-(K-T+\lambda\mathcal{R})h^{ij}+2\lambda\mathcal{R}^{ij}=0,$ (4)
where $T=2\tanh(\rho)-2\lambda\text{sech}^{2}(\rho)$ is a useful parameterization Miao:2023mui . The holographic norm of the displacement operator is given by Miao:2023mui ; Miao:2018dvm
$C_{D}\geq 0$ together with $\lambda\geq 0$ yields the constraints: $0\leq\lambda$ for $\rho\geq 0$; $20\leq\lambda\leq-\frac{\coth(\rho)}{2}\geq 1/$ 2 for $\rho\leq 0$. Note that $\lambda=1/2$ is a critical DGP parameter. Below, we will show $\lambda=1/2$ is a phase-transition point. In the followings, we show only key points. See appendix for detailed derivations.
The vacuum of a strip is dual to AdS soliton in bulk Fujita:2011fp
$ds^{2}=\frac{\frac{dz^{2}}{h(z)}+h(z)d\theta^{2}-dt^{2}+dy^{2}}{z^{2}},$ (6)
where $h(z)=1-z^{3}/z_{h}^{3}$. By using the rescale invariance of AdS, we set $z_{h}=1$. The absence of conical singularities fixes the period of angle $\theta$ in bulk $\beta=4\pi/|h^{\prime}(1)|=4\pi/3$. The strip is defined on the AdS boundary $z=0$ with $0\leq\theta\leq L$, where $L\geq 0$ is the strip width. We derive the energy density $T_{tt}=-1$ from (6) by applying the holographic renormalization deHaro:2000vlm . Compared with the Casimir effect (2), we get the holographic Casimir amplitude
$\kappa_{1}=L^{3},$ (7)
where $L$ can be fixed by NBC (4). For negative brane tension $T\leq 0$, we have
where $H=\frac{1}{2}h(z)\left(1+2\lambda\sqrt{h\left(z_{\max}\right)}+X\right)$, $X=\sqrt{1+4\lambda^{2}h(z)-4\lambda\sqrt{h\left(z_{\max}\right)}}$ and $z_{\text{max}}$ is the turning point obeying $T=-2\sqrt{h(z_{\text{max}})}$. The case with $T>0$ can be obtained from the complement of the first case $L_{\text{II}}=\beta-L_{\text{I}}\Big{(}T\to-T,\lambda\to-\lambda\Big{)}$. This is the typical trick used in Fujita:2011fp , based on extrinsic curvature flip signs when crossing the brane. In the viewpoint of NBC (4), $T,\lambda$ change signs equivalently. Some comments are in order. First, the integrand of (8) cannot take absolute value, which would miss the case of $T<0$ with $\rho>0$ and cannot produce the correct asymptotical behavior, i.e. $\theta=-\sinh(\rho)z+O(z^{2})$ and $(\theta-L)=\sinh(\rho)z+O(z^{2})$. Second, $L_{\text{I}}$ (8) is negative for sufficiently large $\lambda$. Thus, there is an upper bound of $\lambda$. In fact, the constraint $L\geq 0$ yields $\lambda\leq 1$. Third, the case $20\leq\lambda<1/$ 2 can continuously transform into that of Einstein gravity, while the case $1/2<\lambda\leq 1$ cannot. See Fig. 1 for instance. We name them normal and singular phases, respectively. Fourth, the strip width $L$ decreases with brane tension $T$ Fujita:2011fp and $L\geq 0$ sets an upper bound of $T$ Miyaji:2021ktr . For $T\geq 0$, $L$ should be smaller than angle period $\beta$ to avoid conical singularity. Luckly, $L\leq\beta$ is satisfied automatically for $T\geq 0$ Miyaji:2021ktr . On the other hand, there is no upper bound of $L$ for $T<0$, since the conical singularity is hidden behind the brane Tadashi . Thus, it is irrelevant to the bulk dual of strip. As a result, there is no extra constraint on the lower bound of $T$ and it can take the minimal value $T=-2$ ($\rho\to-\infty$) Miyaji:2021ktr ; Tadashi .
We can derive the ratio $(-\kappa_{1}/C_{D})$ from (5,7,8). See Fig. 1 for example. Fig. 1 (up) denotes the normal phase with $20\leq\lambda<1/$ 2, which shows that the smaller $\rho$, the smaller $(-\kappa_{1}/C_{D})$. Besides, all curves approach the same lower bound (purple curve) from the above in the limit $T\to-2$. On the other hand, Fig. 1 (down) labels the singular phase with $1/2\leq\lambda\leq 1$. It shows that ($-\kappa_{1}/C_{D}$) with $\lambda>1/2$ is always larger than that of Einstein gravity (blue curve). Besides, the minimum of ($-\kappa_{1}/C_{D}$) increases with $\lambda$ and approaches zero when $\lambda\to 1$. Since we are interested in the lower bound of ($-\kappa_{1}/C_{D}$), we focus on the normal phase below.
Figure 1: $(-\kappa_{1}/C_{D})$ in normal phase with $20\leq\lambda<1/$ 2 (up) and singular phase with $1/2<\lambda\leq 1$ (down). The up figure shows all curves approach the holographic limit (purple curve) from the above for $\rho\to-\infty$ ($T\to-2$) in the normal phase. The down figure shows $(-\kappa_{1}/C_{D})$ increases with $\lambda$ for $11/2<\lambda<$ 1. In the limit $\lambda\to 1$, we have $\kappa_{1}/C_{D}\to 0$ in the singular phase. $\lambda=1/2$ is the phase transition point.
$L_{I}=\frac{2\sqrt{\pi}\Gamma\left(\frac{4}{3}\right)}{\Gamma\left(\frac{5}{6}\right)}\Big{(}\frac{(1-2\lambda)^{2}}{x}\Big{)}^{\frac{1}{6}}+O\left(\frac{x}{(1-2\lambda)^{2}}\right)^{\frac{5}{6}}$ (9)
$\lim_{T\to-2}(\frac{-\kappa_{1}}{C_{D}})=\lim_{x\to 0}(\frac{-L_{\text{I}}^{3}}{C_{D}})=-\frac{\pi^{5/2}\Gamma\left(\frac{1}{3}\right)^{3}}{108\Gamma\left(\frac{5}{6}\right)^{3}}.$ (10)
It verifies our claim that different holographic models give the same lower bound (3) in the normal phase. Note that (9) implies $L/\beta\to\infty$ for $x\to 0$ ($T\to-2$). Note also that the strip width $L\to Lz_{h}$ can remain finite when recovering $z_{h}\to 0$ in $h(z)$ (6). At the phase transition point $\lambda=1/2$, the lower bound (10) can be achieved by $L=\beta$. See appendix for derivations. The lower bound of $(-\kappa_{1}/C_{D})$ is saturated by negative brane tension. We remark that the negative brane tension is well-defined in AdS/BCFT Fujita:2011fp ; Takayanagi:2011zk ; Miyaji:2021ktr ; Tadashi . First, the brane tension is a cosmological constant rather than a kinetic-energy term on the end-of-the-wolrd (EOW) brane. Of course, the negative cosmological constant is well-defined in AdS/CFT. Second, negative brane tension generally yields a negative A-type boundary central charge. But nothing goes wrong. Recall that the free scalar with the DBC has a negative A-type boundary central charge Jensen:2015swa .
Tests — Let us test the holographic lower bound of the Casimir effect (3) by field theories. First, consider free CFTs. We have $(-\kappa_{1}/C_{D})=-3\pi\zeta(3)/8\approx-1.42$ for free fermions, and $(-\kappa_{1}/C_{D})=-\pi\zeta(3)/2\approx-1.89$ for free scalars with RBC and DBC. Next, consider CFTs with interactions. For Ising model with $\phi^{4}$ interaction at the critical point, we have $2\kappa_{1}\approx 0.820$ and $C_{D}\approx 0.193$ Toldin:2021kun ; Ising referee1 ; Ising referee2 , which yields $(-\kappa_{1}/C_{D})\approx-2.12$. Near the strip boundary, the scalar operator behaves as Cardy:1990xm
$\langle\sigma(z)\rangle=\frac{a_{\sigma}}{(2z)^{\Delta_{\sigma}}}\Big{(}1+B_{\sigma}(\frac{z}{L})^{d}+...\Big{)},$ (11)
@@
Here, we use the notation of Toldin:2021kun with $d=3$ for 3d CFTs. For $O(N)$ model with $N=2,3$, Toldin:2021kun obtains $B_{\sigma}\approx 1.21,1.07$ and $\Delta_{\sigma}\approx 0.519088,0.518920$ by applying Monte Carlo simulations. Then (12) gives $(-\kappa_{1}/C_{D})\approx-1.83,-1.62$ for $N=2,3$, respectively. We summarize all the above results in Table.1, which all obey the holographic bound (3). It strongly supports our proposal. Interestingly, (12) suggests the bound of the Casimir effect imposes an upper bound of $B_{\sigma}/\Delta_{\sigma}$. Below we briefly comment on various generalizations of our proposal (3). We show only key points and leave detailed discussions and open questions to future work.
@@
General boundaries — Holography is expected to set a bound of Casimir effect for general boundary shapes, not be limited to strips. Let us provide some evidence. We first consider the wedge space, which is the simplest generalization of a strip. The Casimir effect takes the form
@@
where $f(\Omega)$ is the Casimir amplitude and $\Omega$ is the opening angle of wedge. In the limit $\Omega\to 0,r\to\infty$ with $r\Omega=L$ fixed, the large $r$ region of the wedge approaches a strip. It results in $f(\Omega)\to\kappa_{1}/\Omega^{3}$ for $\Omega\to 0$ Deutsch:1978sc . On the other hand, we have $f(\Omega)\to(C_{D}\pi/64)(\pi-\Omega)$ in the smooth limit $\Omega\to\pi$ Miao:2024ddp . These two limits suggest
$(\frac{-f(\Omega)}{C_{D}})\geq(\frac{-f(\Omega)}{C_{D}})_{\text{holo}},\ \text{for}\ 0<\Omega\leq\pi,$ (14)
where the holographic bound is derived in the limit $T\to-2$. By applying the method of Miao:2024ddp and this letter, we numerically verify (14) holds. Let us go on to discuss the general boundary. Near any smooth boundary, the normal-normal component of stress tensor takes a universal form in flat space Deutsch:1978sc ; Miao:2017aba
@@
$(\frac{\bar{t}_{nn}}{C_{D}})\geq(\frac{\bar{t}_{nn}}{C_{D}})_{\text{holo}},$ (16)

### v3

@@ Abstract
Inspired by the Kovtun-Son-Starinet bound, we propose that holography imposes a lower bound on the Casimir effect. For simplicity, we focus on the Casimir effect between parallel planes for three-dimensional conformal field theories and briefly comment on the generalizations to other boundary shapes and higher dimensions. Remarkably, the ghost-free holographic models impose a universal lower bound on the Casimir effect. We verify the holographic bound by free theories, the Ising model, and $O(N)$ models with $N=2,3$ at critical points and prove it for the two-dimensional case. Remarkably, a general class of quantum field theories without conformal symmetries also obeys the holographic bound.
Introduction — Casimir effect Casimir:1948dh is an interesting quantum effect that originates from the changes of vacuum fluctuations due to the boundary Plunien:1986ca ; Bordag:2001qi ; Milton:2004ya ; Bordag:2009zz . It is a candidate of dark energy Wang:2016och and plays a vital role in QCD Vepstas:1984sw and wormhole physics Morris:1988tu ; Maldacena:2020sxe . It has been measured in experiments Mohideen:1998iz ; Bressi:2002fr ; Klimchitskaya:2009cw and has significant potential applications in nanotechnology. As the vacuum energy, Casimir energy is defined as the lowest energy of a quantum system with a boundary. Since energy should be bounded from below, it is expected that there is a fundamental lower bound of the Casimir effect. This letter explores this vital problem. Holography (AdS/CFT) Maldacena:1997re provides a powerful tool to study the strongly coupled conformal field theory (CFT) and usually imposes bounds on various physical phenomena. The most famous example is the Kovtun-Son-Starinet (KSS) bound Kovtun:2004de for hydromechanics, which conjectures that the ratio of shear viscosity $\eta$ to entropy density $s$ has a lower bound
@@
where $\hbar$ is the reduced Planck constant and $k_{B}$ is the Boltzmann constant. The lower bound is obtained from the strongly coupled CFT dual to Einstein gravity Policastro:2001yc and can be modified slightly by higher derivative gravity Brigante:2007nu ; Kats:2007mq ; Brigante:2008gz .
Now, let us explore whether AdS/CFT sets a similar lower bound for the Casimir effect. For simplicity, we focus on the Casimir effect for three-dimensional (3D) CFTs between parallel planes (strip) and comment on various generalizations at the end of the Letter. We consider flat space and impose the same boundary conditions on the two planes. Then, the Casimir effect takes a universal form Miao:2024ddp
@@
where $\kappa_{1}$ is the dimensionless Casimir amplitude, $c$ is the speed of light, and $L$ is the strip width. For the fixed $L$, $\kappa_{1}$ should be bounded from below to avoid negative infinity energy density $T_{tt}=-\kappa_{1}\hbar c/L^{3}$ and pressure $T_{nn}=-2\kappa_{1}\hbar c/L^{3}$. Note that the Casimir amplitude $\kappa_{1}$ is an extensive parameter in the sense that if one CFT has a Casimir amplitude $\kappa_{1}$, then $N$-copies of the same CFT have a Casimir amplitude $N\kappa_{1}$. It implies that only $\kappa_{1}$ over another extensive parameter could have a lower bound. Natural candidates of these extensive parameters are boundary central charges in Weyl anomalyJensen:2015swa $\mathcal{A}=\int_{\partial M}d^{2}x\sqrt{|\sigma|}\Big(aR_{\partial M}+b\text{tr}\bar{k}^{2}\Big)$, where $R_{\partial M}$ and $\bar{k}$ are the intrinsic Ricci scalar and traceless parts of extrinsic curvatures on the boundary $\partial M$, respectively. $a$ is the A-type boundary central charge, which decreases under the boundary renomalization group (RG) flow Jensen:2015swa . We remark that $a$ can be positive, zero, or negative. For example, $a=\pm 1/(384\pi)$ for conformally coupled free scalar with Robin boundary condition (RBC) and Dirichlet boundary condition (DBC) Jensen:2015swa , respectively; $a\sim\sinh(\rho)$ with $\infty<\rho<\infty$ for holographic CFT3 Fujita:2011fp . Since $a$ can be zero and negative, there is no lower bound for $\kappa_{1}/a$. On the other hand, the B-type boundary central charge $b$ is always positive since it is related to the Zamolodchikov norm of the displacement operator, i.e., $b=C_{D}\pi/32>0$. The displacement operator $D(y)$ describes the violation of space translation invariance normal to the boundary Billo:2016cpy , i.e., $\nabla_{i}T^{ij}=\delta(x)D(y)n^{j}$. The two-point function defines its Zamolodchikov norm, i.e., $\langle D(y)D(0)\rangle=C_{D}c^{2}\hbar^{2}/|y|^{6}$. By definition, $C_{D}$ is dimensionless and positive. The above arguments suggest $\kappa_{1}/C_{D}$ should have a physical lower bound. Usually, a bound is set either by free theories or strongly coupled CFTs dual to gravity. We have checked all the known examples and find that, similar to the KSS bound, holography gives the smallest $\kappa_{1}/C_{D}$. Thus, we propose that the ratio of Casimir amplitude to the norm of the displacement operator has a lower bound set by holography
$-\frac{\kappa_{1}}{C_{D}}\geq-\frac{\pi^{5/2}\Gamma\left(\frac{1}{3}\right)^{3}}{108\Gamma\left(\frac{5}{6}\right)^{3}}\approx-2.17,\ \ \ \text{for 3D CFTs}.$ (3)
It means the energy density $\kappa_{1}/L^{3}$ per degree of freedom $C_{D}$ is bounded from below for a physical system with fixed size $L$. Remarkably, unlike the KSS bound (1), different holographic models give the same universal lower bound (3) for the Casimir effect. We verify the lower bound (3) by free CFTs, the Ising model, and the $O(N)$ model with $N=2,3$ at critical points. Furthermore, we prove the corresponding bound for two-dimensional (2D) CFTs, which strongly supports our general proposal. Below, we discuss the test of our proposal, the derivation of the holographic bound, and its various generalizations. For simplicity, we focus on natural units with $c=\hbar=k_{B}=1$ in the following.
Table 1: $(-\kappa_{1}/C_{D})$ for various 3D CFTs
Dirac fermion Scalar Ising O(2) O(3) AdS/BCFT
1.42-1.42 1.89-1.89 -2.12 -1.83 -1.62 2.17-2.17
Tests — Let us test the holographic bound (3) by various 3D CFTs, which is summarized in Table 1. We have $\kappa_{1}/C_{D}=-3\pi\zeta(3)/8\approx-1.42$ for Dirac fermions with bag boundary condition (BC) Bellucci:2009hh ; McAvity:1993ue , and $\kappa_{1}/C_{D}=-\pi\zeta(3)/2\approx-1.89$ for free scalars with RBC and DBC Romeo:2000wt ; Miao:2018dvm . For Ising model with $\phi^{4}$ interaction at the critical point, we have $2\kappa_{1}\approx 0.820$ and $C_{D}\approx 0.193$ Toldin:2021kun ; Ising referee1 ; Ising referee2 , which yields $\kappa_{1}/C_{D}\approx-2.12$. Here, we impose with $(++)$ BC for the Ising model, where the spins on the two boundaries have the same signs. Near the strip boundary, the scalar operator behaves as Cardy:1990xm
@@
Here, we use the notation of Toldin:2021kun with $d=3$ for 3D CFTs. For $O(N)$ model with $N=2,3$ and open BCs, Toldin:2021kun obtains $B_{\sigma}\approx 1.21,1.07$ and $\Delta_{\sigma}\approx 0.519088,0.518920$ by applying Monte Carlo simulations. Then (5) gives $\kappa_{1}/C_{D}\approx-1.83,-1.62$ for $N=2,3$, respectively. We summarize all the above results in the Table 1, which all obey the holographic bound (3) will be derived below. It strongly supports our proposal. Interestingly, (5) suggests the bound of the Casimir effect imposes an upper bound of $B_{\sigma}/\Delta_{\sigma}$.
Holographic bound — Let us study the holographic bound of the Casimir effect of a strip in AdS/BCFT Takayanagi:2011zk . We are interested in the general ghost-free gravity in AdS4, which turns out to be Dvali-Gabadadze-Porrati (DGP) gravity,Dvali:2000hr
$I=\int_{N}d^{4}x\sqrt{|g|}\Big(R+6\Big)+2\int_{Q}d^{3}y\sqrt{|h|}\Big(K-T+\lambda\mathcal{R}\Big),$ (6)
where $K$ is the extrinsic curvature, $T$ is the brane tension, $\lambda$ is the DGP parameter, $R$ and $\mathcal{R}$ are Ricci scalars in bulk $N$ and on the brane $Q$, respectively. We have set Newton’s constant $16\pi G_{N}=1$ and AdS radius $l=1$ for simplicity. Note that higher derivative gravity and negative DGP gravity generally suffer the ghost problems Miao:2023mui , while Gauss-Bonnet/Lovelock gravity appears only in dimensions higher than four. Thus, positive DGP gravity with $\lambda\geq 0$ is the general ghost-free gravity theory in AdS4. The case $\lambda=0$ corresponds to Einstein gravity. Following Takayanagi:2011zk , we impose the Neumann boundary condition (NBC) on the brane $Q$
$K^{ij}-(K-T+\lambda\mathcal{R})h^{ij}+2\lambda\mathcal{R}^{ij}=0,$ (7)
where $T=2\tanh(\rho)-2\lambda\text{sech}^{2}(\rho)$ is a useful parametrization Miao:2023mui . The holographic norm of the displacement operator is given by Miao:2023mui ; Miao:2018dvm
$C_{D}\geq 0$ together with $\lambda\geq 0$ yields the constraints: $0\leq\lambda$ for $\rho\geq 0$; $0\leq\lambda\leq-\frac{\coth(\rho)}{2}\geq 1/2$ for $\rho\leq 0$. Below, we will show that $\lambda=1/2$ is a phase-transition point and there are other physical constraints for $\lambda$. In the following, we show only key points and leave the detailed derivations to the supplemental material.
The vacuum of a strip is dual to the AdS soliton Fujita:2011fp
$ds^{2}=\frac{\frac{dz^{2}}{h(z)}+h(z)d\theta^{2}-dt^{2}+dy^{2}}{z^{2}},$ (9)
where $h(z)=1-z^{3}/z_{h}^{3}$. The absence of conical singularities fixes the period of angle $\theta$ in bulk $\beta=4\pi/|h^{\prime}(z_{h})|=4z_{h}\pi/3$. The strip is defined on the AdS boundary $z=0$ with $0\leq\theta\leq L$, where $L\geq 0$ is the strip width. We derive the energy density $T_{tt}=-1/z_{h}^{3}$ from (9) by applying the holographic renormalization deHaro:2000vlm . Compared with the Casimir effect (2), we get the holographic Casimir amplitude
$\kappa_{1}=L^{3}/z_{h}^{3},$ (10)
where $L$ can be fixed by NBC (7). By using the rescale invariance of AdS, we set $z_{h}=1$ below. For negative brane tension $T\leq 0$, we derive the strip width
where $H=\frac{1}{2}h(z)\left(1+2\lambda\sqrt{h\left(z_{\max}\right)}+X\right)$, $X=\sqrt{1+4\lambda^{2}h(z)-4\lambda\sqrt{h\left(z_{\max}\right)}}$ and $z_{\text{max}}$ is the turning point obeying $T=-2\sqrt{h(z_{\text{max}})}$. The strip width with $T>0$ can be obtained from the complement of that with $T<0$ by $L_{\text{II}}=\beta-L_{\text{I}}\Big(T\to-T,\lambda\to-\lambda\Big)$. It is the typical trick used in Fujita:2011fp , based on extrinsic curvature flip signs when crossing the brane. In the viewpoint of NBC (7), $T,\lambda$ change signs equivalently. Some comments are in order. First, $L_{\text{I}}$ (11) is negative for sufficiently large $\lambda$. The physical constraint $L\geq 0$ yields $\lambda\leq 1$. Second, the case $0\leq\lambda<1/2$ can continuously transform into that of Einstein gravity, while the case $1/2<\lambda\leq 1$ cannot. See Fig. 1 for instance. We name them normal and singular phases, respectively. Third, the strip width $L$ decreases with brane tension $T$ Fujita:2011fp and $L\geq 0$ sets an upper bound of $T$ Miyaji:2021ktr . For $T\geq 0$, $L$ should be smaller than the angle period $\beta$ to avoid conical singularity. Luckily, $L\leq\beta$ is satisfied automatically for $T\geq 0$ Miyaji:2021ktr . On the other hand, there is no upper bound of $L$ for $T<0$, since the conical singularity is hidden behind the brane Tadashi . Thus, it is irrelevant to the bulk dual of the strip. As a result, there is no extra constraint on the lower bound of $T$, and it can take the minimal value T=2T=-2 Miyaji:2021ktr ; Tadashi . We remark that $T\geq-2$ sets no constraint on $\rho$ in the normal phase, while it sets a lower bound of $\rho$ in the singular phase. See Fig. 1 (bottom).
We can derive the ratio $\kappa_{1}/C_{D}$ from (8,10,11). Refer to Fig. 1 for details. Fig. 1 (top) denotes the normal phase with $0\leq\lambda<1/2$, which shows that the smaller $\rho$, the smaller the $\kappa_{1}/C_{D}$. Besides, all curves approach the same lower bound (red line) from above in the limit $T\to-2$. On the other hand, Fig. 1 (bottom) labels the singular phase with $1/2\leq\lambda\leq 1$. It shows that $\kappa_{1}/C_{D}$ with $\lambda>1/2$ is always larger than that of Einstein gravity (blue curve). Besides, the minimum of $\kappa_{1}/C_{D}$ increases with $\lambda$ and approaches zero when $\lambda\to 1$. Since we are interested in the lower bound of ($\kappa_{1}/C_{D}$), we focus on the normal phase below.
Let us explain the physical reasons why the minimal brane tension yields the lower bound of the Casimir effect in the normal phase. First, the Casimir energy equals the free energy at zero temperature, which can be calculated holographically by the Euclidean gravitational action. Naturally, the smaller the brane tension (energy), the smaller the Euclidean action, and, thus, the smaller the Casimir energy. Second, the brane tension decreases under boundary RG flow, i.e., $T_{\text{IR}}<T_{\text{UV}}$ Fujita:2011fp . At IR, many modes become inactive, resulting in reduced energy. Third, the Casimir effect is driven by long-range quantum correlations Dantchev:2022hvy . As the tension (energy) decreases, these quantum correlations become stronger, which in turn increases the Casimir amplitude and reduces the Casimir energy. In summary, the minimal brane tension corresponds to the minimal Casimir energy.
Figure 1: $\kappa_{1}/C_{D}$ in normal phase with $0\leq\lambda<1/2$ (top) and singular phase with $1/2<\lambda\leq 1$ (bottom) for DGP gravity. Einstein gravity corresponds to $\lambda=0$. The top figure shows all holographic curves approach the same limit (red line) from above for $\rho\to$-$\infty$ ($T\to-2$) in the normal phase. The bottom figure shows $\kappa_{1}/C_{D}$ increases with $\lambda$ and approaches zero for $\lambda\to 1$ in the singular phase. Together, it shows all examples obey the holographic bound eq.(3) (red line).
$L_{\text{I}}=\frac{2\sqrt{\pi}\Gamma\left(\frac{4}{3}\right)}{\Gamma\left(\frac{5}{6}\right)}\Big(\frac{(1-2\lambda)^{2}}{x}\Big)^{\frac{1}{6}}+O\left(\frac{x}{(1-2\lambda)^{2}}\right)^{\frac{5}{6}}.$ (12)
$\lim_{T\to-2}\frac{-\kappa_{1}}{C_{D}}=\lim_{x\to 0}\frac{-L_{\text{I}}^{3}}{C_{D}}=-\frac{\pi^{5/2}\Gamma\left(\frac{1}{3}\right)^{3}}{108\Gamma\left(\frac{5}{6}\right)^{3}}.$ (13)
It verifies our claim that different holographic models give the same lower bound (3) in the normal phase. The lower bound of $(-\kappa_{1}/C_{D})$ is saturated by negative brane tension. We remark that the negative brane tension is well defined in AdS/BCFT Fujita:2011fp ; Takayanagi:2011zk ; Miyaji:2021ktr ; Tadashi . First, the brane tension is a cosmological constant rather than a kinetic-energy term on the end-of-the-world brane. Of course, the negative cosmological constant is well defined in AdS/CFT. Second, negative brane tension generally yields a negative A-type boundary central charge. But nothing goes wrong. Recall that the free scalar with the DBC has a negative A-type boundary central charge Jensen:2015swa .
@@
General boundaries — Holography is expected to set a bound on the Casimir effect for general boundary shapes, not limited to strips. Let us provide some evidence. We first consider the wedge space, which is the simplest generalization of a strip. The Casimir effect takes the form
@@
where $f(\Omega)$ is the Casimir amplitude and $\Omega$ is the opening angle of wedge. In the limit $\Omega\to 0,r\to\infty$ with $r\Omega=L$ fixed, the large $r$ region of the wedge approaches a strip. It results in $f(\Omega)\to\kappa_{1}/\Omega^{3}$ for $\Omega\to 0$ Deutsch:1978sc . On the other hand, we have $f(\Omega)\to(C_{D}\pi/64)(\pi-\Omega)$ in the smooth limit $\Omega\to\pi$ Miao:2024ddp . These two limits suggest AdS/BCFT with minimal brane tension sets the lower bound of the wedge Casimir effect
$\frac{-f(\Omega)}{C_{D}}\geq\lim_{T\to-2}(\frac{-f(\Omega)}{C_{D}})_{\text{holo}},\ \text{for}\ 0<\Omega\leq\pi.$ (15)
See the following long paper Miao:2025utb for the derivation and test of (15). Let us go on to discuss the general boundary. Near any smooth boundary, the normal-normal component of the stress tensor takes a universal form in flat space Deutsch:1978sc ; Miao:2017aba
@@
$(\frac{\bar{t}_{nn}}{C_{D}})\geq\lim_{T\to-2}(\frac{\bar{t}_{nn}}{C_{D}})_{\text{holo}},$ (17)
where $\bar{t}_{nn}=(\int_{\partial M}t_{nn}dS)/(\int_{\partial M}dS)$ is related to the variation of the total energy concerning the system size, i.e., dE/dL-dE/dL.
General dimensions — Let us comment on the holographic bound of the Casimir effect for a strip in general dimensions. The corresponding Casimir amplitude $\kappa_{1}$ and norm of displacement operator $C_{D}$ are given by Miao:2024ddp for Einstein gravity. By taking the limit $T\to-(d-1)$, we derive a universal lower bound,
$\frac{-\kappa_{1}}{C_{D}}\geq\frac{-2^{d-2}d\pi^{d-\frac{1}{2}}\Gamma\left(\frac{d-1}{2}\right)\Gamma\left(\frac{1}{d}\right)^{d}}{\Gamma(d+2)\left(d\Gamma\left(\frac{1}{2}+\frac{1}{d}\right)\right)^{d}}.$ (18)